# Linear fractional relative risk aversion

### v1

@@ 3.1 Price equilibrium
Second, assume that $\sigma=1$. In this case, $\beta\neq 1$ must hold because of the assumption that $1-\beta\sigma>0$. Thus, $\Phi(q)>0$ implies $q<-\frac{\beta}{2\alpha}$. We also know from Table 3 that $0\leq q<-\frac{\beta}{\alpha}$ must hold. Since $\frac{\beta}{2\alpha}<-\frac{\beta}{\alpha}$ holds, the second-order condition holds for all $0\leq q<-\frac{\beta}{2\alpha}$ but it does not hold for $\frac{\beta}{2\alpha}\leq q<-\frac{\beta}{\alpha}$. Since the range of $q$ is modified, we need to modify that for $\mu$. Noting that $\frac{\partial\mu}{\partial q}=\frac{\partial}{\partial q}\bigl{(}1-\frac{\alpha q+\beta}{\alpha\sigma q+1}\bigr{)}=-\frac{\alpha(1-\beta\sigma)}{(\alpha\sigma q+1)^{2}}>0$, we evaluate $\mu=1-\frac{\alpha q+\beta}{\alpha\sigma q+1}$ at $q=0$ and $q=-\frac{\beta}{2\alpha}$ to obtain $1-\beta\leq\mu<\frac{2(1-\beta)}{2-\beta}<1$. Hence, we obtain the range of markups $1<\frac{2-\beta}{2(1-\beta)}<\frac{1}{\mu}\leq\frac{1}{1-\beta}$ for $\sigma=1$.
@@
Last, assume that $\sigma>1$. In this case, $\beta\neq 1$ must hold because of the assumption that $1-\beta\sigma>0$. Thus, $\Phi(q)=0$ has two real roots (26). Both roots are positive since $1-\beta+\Delta>0$ and $1-\beta-\Delta>0$. Since $\Phi$ is convex, it is negative between its roots. Since $\frac{1-\beta-\Delta}{\alpha(\sigma-1)}<-\frac{\beta}{\alpha}<-\frac{1-\beta+\Delta}{\alpha(\sigma-1)}$, the second-order condition holds for all $0\leq q<-\frac{1-\beta-\Delta}{\alpha(\sigma-1)}$ but does not hold for $\frac{1-\beta-\Delta}{\alpha(\sigma-1)}\leq q<-\frac{\beta}{\alpha}$. Since the bounds for $q$ are modified, we need to modify those for $\mu$. Noting that $\frac{\partial\mu}{\partial q}=\frac{\partial}{\partial q}\bigl{(}1-\frac{\alpha q+\beta}{\alpha\sigma q+1}\bigr{)}=-\frac{\alpha(1-\beta\sigma)}{(\alpha\sigma q+1)^{2}}>0$, we evaluate $\mu=1-\frac{\alpha q+\beta}{\alpha\sigma q+1}$ at $q=0$ and $q=-\frac{1-\beta-\Delta}{\alpha(\sigma-1)}<-\frac{\beta}{\alpha}$ to obtain $1-\beta\leq\mu<\frac{(1-\beta)\sigma+\Delta}{1+(1-\beta)\sigma}<1$. Hence, we obtain the bounds for markups $1<\frac{1+(1-\beta)\sigma}{(1-\beta)\sigma+\Delta}<\frac{1}{\mu}\leq\frac{1}{1-\beta}$ for $\sigma>1$.
@@
Since $\beta\neq 1$ and $1-\beta\sigma<0$, $\Phi(q)=0$ has no real roots. Since $\sigma>1$ and $\beta\not\in\{0,1\}$, $\Phi$ is convex and $\Phi(0)=\beta(1-\beta)>0$, which implies that $\Phi$ is positive for all $q\geq 0$. Hence, the SOC holds for all $q\geq 0$. Therefore, the SOC holds for $0\leq q<-\frac{1-\beta}{\alpha(\sigma-1)}$ and thus for $\frac{1}{1-\beta}\leq\frac{1}{\mu}<\infty$ given in Table 3 when $\beta\in(0,1)$.
Finally, observe from Table 4 that marginal cost pricing $\frac{1}{\mu}=1$ is admissible only when $\beta=0$ in Case 1a. Note that $\frac{\sigma}{\sigma-1}>1$ holds in Case 1b, $\frac{(1-\beta)\sigma+1}{(1-\beta)\sigma+\Delta}>1$ holds in Case 2a, and that $\beta=0$ is not admissible in Case 2b.
Case 1: $\alpha>0$.
Markups Quantities
$\sigma\leq 1$ $\sigma>1$ $\sigma\leq 1$ $\sigma>1$
Case 2: $\alpha<0$.
Markups Quantities
$\sigma\leq 1$ $\sigma>1$ $\sigma\leq 1$ $\sigma>1$
Using the marginal utility function (9), which nests the HARA ($\beta=0$) and CREMR ($\beta=1$) cases in (16), the first-order condition for utility maximization is given by $u^{\prime}(q)=K(1+\alpha\sigma q)^{\beta-\frac{1}{\sigma}}q^{-\beta}=\nu p$. We can rewrite it as $(1+\alpha\sigma q)^{\beta-\frac{1}{\sigma}}q^{-\beta}=\frac{\nu p}{K}=\frac{\nu m}{K}\frac{p}{m}$. Letting $x\equiv\frac{\nu m}{K}>0$ and $\mu=\frac{m}{p}$, we have
$(1+\alpha\sigma q)^{\beta-\frac{1}{\sigma}}q^{-\beta}=\frac{x}{\mu}.$ (27)
Proposition 4 (Existence and uniqueness of a markup function)
If $\alpha\neq 0$ and $\beta\neq\frac{1}{\sigma}$, then there exists a unique markup function $\frac{1}{\mu}=\frac{1}{{\cal W}(x,\alpha,\beta,\sigma)}\geq 1$ satisfying the first-order condition in Table 5 and the second-order condition (24) conditional on the restrictions on $x$ in Table 6. If $\alpha=0$ and $\beta\in(0,1)$, then there exists a unique markup function $\frac{1}{\mu}=\frac{1}{{\cal W}(x,0,\beta,\sigma)}=\frac{1}{1-\beta}>1$ satisfying the first-order condition (21) and the second-order condition (24). If $\beta=\frac{1}{\sigma}$ and $\sigma>1$, then there exists a unique markup function $\frac{1}{\mu}=\frac{1}{{\cal W}(x,\alpha,\frac{1}{\sigma},\sigma)}=\frac{\sigma}{\sigma-1}>1$ satisfying the first-order condition (21) and the second-order condition (24). [^28] Alternatively, assume that $\sigma>1$. Then, recalling $x=\frac{\nu m}{K}$, using $K=(1-\beta)(1+\alpha\sigma)^{\frac{1}{\sigma}}+\beta(1+\alpha\sigma)^{\frac{1}{\sigma}-1}$, and letting $\alpha\to\infty$ in the bottom-right expression of Table 5, we can show that there exists a unique markup function $\frac{1}{\mu}=\lim_{\alpha\to\infty}\frac{1}{{\cal W}(x(\alpha),\alpha,\beta,\sigma)}=\frac{\sigma}{\sigma-1}>1$. Furthermore, when $\alpha\to\infty$, the second-order condition (24) reduces to $\frac{\sigma-1}{\sigma}>0$, which holds for all $\sigma>1$.
$\beta=0$ $\beta\in(0,1]$
Proof.
@@
$\mu\left[\frac{1-\beta\sigma}{1-\sigma(1-\mu)}\right]^{\beta-\frac{1}{\sigma}}\left[\frac{1}{\alpha}\frac{1-\beta-\mu}{1-\sigma(1-\mu)}\right]^{-\beta}=x.$ (28)
Consider the subcase with $\beta\in(0,1]$. To obtain the bottom-right expression of Table 5 from (28), we first assume that $\sigma\neq 0$. In that case, we can rewrite (28) as follows:
@@
When $\sigma$ goes to 0, equation (28) reduces to $\mu{\rm e}^{-(1-\beta-\mu)}\left(\frac{1-\beta-\mu}{\alpha}\right)^{-\beta}=x,$ which can be rewritten as
@@
Consider the subcase with $\beta=0$. Then, (28) implies $\mu[\frac{1}{1-\sigma(1-\mu)}]^{-\frac{1}{\sigma}}=x$. We can rewrite it as $\frac{1}{1-\sigma(1-\mu)}=(\frac{x}{\mu})^{-\sigma}$, so that $1-\sigma(1-\mu)=(\frac{\mu}{x})^{-\sigma}$. Hence, we obtain $1-\mu=\frac{1-(\frac{\mu}{x})^{-\sigma}}{\sigma}$. When $\sigma$ goes to 0, the foregoing equation reduces to $1-\mu=\ln(\frac{\mu}{x})$.
@@
We now prove the existence and uniqueness of a markup function. The right-hand side of expression (28) is exogenous for each firm, whereas the left-hand side depends on the markup $\mu$ that is endogenously determined. Differentiating the left-hand side of (28) with respect to $\mu$ yields
$\frac{\partial{\rm LHS}}{\partial\mu}=\frac{1-\beta-\mu^{2}-\sigma(1-\beta)(1-\mu)^{2}}{(1-\beta-\mu)[1-\sigma(1-\mu)]}\frac{x}{\mu}.$ (29)
@@
$\frac{1-\beta-\mu^{2}-\sigma(1-\beta)(1-\mu)^{2}}{(1-\mu)(1-\beta\sigma)}>0,$ (30)
@@
Consider the subcase with $\beta\in(0,1)$. In this case, $q>0$ must hold as explained before, so that $1-\beta-\mu>0$. Since $1-\mu>\beta>0$ and $1-\beta\sigma>0$ hold, the denominator of (30) is positive. We thus obtain $1-\beta-\mu^{2}-\sigma(1-\beta)(1-\mu)^{2}>0$ from (30).
Since $\frac{x}{\mu}>0$, $1-\beta-\mu>0$, and $1-\sigma(1-\mu)>0$, we know that (29) is positive, i.e., the LHS of (28) is strictly increasing in $\mu$. Thus, if a solution $\mu$ to (28) exists, it is unique. The solution exists when the LHS of (28) is smaller than the right-hand side, $x$, as $\mu$ goes to 0 and is greater than $x$ as $\mu$ goes to $1-\beta$. Computing the limits of the LHS, we have
@@
Thus, when $\sigma\leq 1$ and $0<\beta<1$ hold, there exists a unique solution $\mu\in(0,1-\beta)$ to (28) regardless of the value of $x$.
Consider the subcase with $\beta=0$. Setting $\beta=0$ and differentiating the LHS of (28) yield $1-\sigma(1-\mu)]^{-\frac{\sigma-1}{\sigma}}[1-\sigma(1-\mu)+\mu]>0$. Computing the limits of the LHS, we have
@@
Thus, when $\sigma\leq 1$ and $\beta=0$ hold, there exists a unique solution $\mu\in(0,1]$ to (28) only for $x\leq 1$.
@@
Consider the subcase with $\beta\in(0,1)$. In this case, $q>0$ must hold, so that $1-\beta-\mu>0$. Since $1-\mu>\beta>0$ and $1-\beta\sigma>0$ hold, the denominator of (30) is positive. We thus obtain $1-\beta-\mu^{2}-\sigma(1-\beta)(1-\mu)^{2}>0$ from (30).
Since $\frac{x}{\mu}>0$, $1-\beta-\mu>0$, and $1-\sigma(1-\mu)>0$, we know that (29) is positive, i.e., the LHS of (28) is strictly increasing in $\mu$. Thus, if a solution $\mu$ to (28) exists, it is unique. The solution exists when the LHS of (28) is smaller than the right-hand side, $x$, as $\mu$ goes to $\frac{\sigma-1}{\sigma}$ and is greater than $x$ as $\mu$ goes to $1-\beta$. Computing the limits of the LHS, we have
@@
Thus, when $\sigma>1$ and $0<\beta<1$ hold, there exists a unique solution $\mu\in\bigl{(}\frac{\sigma-1}{\sigma},1-\beta\bigr{)}$ to (28) regardless of the value of $x$.
Consider the subcase with $\beta=0$. Setting $\beta=0$ and differentiating the LHS of (28) yield $1-\sigma(1-\mu)]^{-\frac{\sigma-1}{\sigma}}[1-\sigma(1-\mu)+\mu]>0$. Computing the limits of the LHS, we have
@@
Consider the subcase with $\beta\in(0,1)$. The range of markups $\frac{\sigma}{\sigma-1}<\frac{1}{\mu}\leq\frac{1}{1-\beta}$ in Table 4 implies that the range of $\mu$ is given by $1-\beta\leq\mu<\frac{\sigma-1}{\sigma}$. Since $\mu<\frac{\sigma-1}{\sigma}<1$ and $1-\beta\sigma<0$ hold, the denominator of (30) is negative. We thus obtain $1-\beta-\mu^{2}-\sigma(1-\beta)(1-\mu)^{2}<0$ from (30).
Recall that when $\beta\in(0,1)$, $q>0$ must hold, so that $1-\beta-\mu<0$. Since $\frac{x}{\mu}>0$, $1-\beta-\mu<0$, and $1-\sigma(1-\mu)<0$, we know that (29) is negative, i.e., the LHS of (28) is strictly decreasing in $\mu$. Thus, if a solution $\mu$ to (28) exists, it is unique. The solution exists when the LHS of (28) is greater than the right-hand side, $x$, as $\mu$ goes to $1-\beta$ and is smaller than $x$ as $\mu$ goes to $\frac{\sigma-1}{\sigma}$. Computing the limits of the LHS, we have
@@
Thus, when $0<\beta<1$, there exists a unique solution $\mu\in\bigl{(}1-\beta,\frac{\sigma-1}{\sigma}\bigr{)}$ to (28) regardless of the value of $x$.
Consider the subcase with $\beta=1$. The range of markups $\frac{\sigma}{\sigma-1}<\frac{1}{\mu}<\infty$ in Table 4 implies that the range of $\mu$ is given by $0<\mu<\frac{\sigma-1}{\sigma}$. Since $\mu<\frac{\sigma-1}{\sigma}<1$ and $1-\beta\sigma<0$ hold, the denominator of (30) is negative. We thus obtain $1-\beta-\mu^{2}-\sigma(1-\beta)(1-\mu)^{2}<0$ from (30).
Recall that when $\beta=1$, $q>0$ must hold, so that $1-\beta-\mu<0$. Since $\frac{x}{\mu}>0$, $1-\beta-\mu<0$, and $1-\sigma(1-\mu)<0$, we know that (29) is negative, i.e., the LHS of (28) is strictly decreasing in $\mu$. Thus, if a solution $\mu$ to (28) exists, it is unique. The solution exists when the LHS of (28) is greater than the right-hand side, $x$, as $\mu$ goes to 0 and is smaller than $x$ as $\mu$ goes to $\frac{\sigma-1}{\sigma}$. Computing the limits of the LHS, we have
@@
In this case, $\beta=0$ is excluded by Step 1 of the proof of Proposition 3 and $\beta=1$ is not admissible by Step 2 of the proof of Proposition 3. The range of markups $\frac{(1-\beta)\sigma+1}{(1-\beta)\sigma+\Delta}<\frac{1}{\mu}\leq\frac{1}{1-\beta}$ in Table 4 implies that the range of $\mu$ is given by $1-\beta\leq\mu<\frac{(1-\beta)\sigma+\Delta}{(1-\beta)\sigma+1}$. Since $\mu<\frac{(1-\beta)\sigma+\Delta}{(1-\beta)\sigma+1}<1$ and $1-\beta\sigma>0$ hold, the denominator of (30) is positive. We thus obtain $1-\beta-\mu^{2}-\sigma(1-\beta)(1-\mu)^{2}>0$ from (30).
Recall that when $\beta\in(0,1)$, $q>0$ must hold, so that $1-\beta-\mu<0$. Since $\frac{x}{\mu}>0$, $1-\beta-\mu<0$, and $1-\sigma(1-\mu)>0$, we know that (29) is negative, i.e., the LHS of (28) is strictly decreasing in $\mu$. Thus, if a solution $\mu$ to (28) exists, it is unique. The solution exists when the LHS of (28) is greater than the right-hand side, $x$, as $\mu$ goes to $1-\beta$ and is smaller than $x$ as $\mu$ goes to $\overline{\mu}\equiv\frac{(1-\beta)\sigma+\Delta}{(1-\beta)\sigma+1}$. Computing the limits of the LHS, we have
@@
Thus, there exists a unique solution $\mu\in(1-\beta,\overline{\mu})$ to (28) for $x>\underline{x}$.
@@
In this case, $\beta=0$ and $\beta=1$ are not admissible, whereas $\sigma>1$ must hold. The range of markups $\frac{1}{1-\beta}\leq\frac{1}{\mu}<\infty$ in Table 4 implies that the range of $\mu$ is given by $0<\mu\leq 1-\beta$. Since $\mu\leq 1-\beta<1$ and $1-\beta\sigma<0$ hold, the denominator of (30) is negative. We thus obtain $1-\beta-\mu^{2}-\sigma(1-\beta)(1-\mu)^{2}<0$ from (30).
Recall that when $\beta\in(0,1)$, $q>0$ must hold, so that $1-\beta-\mu>0$. Since $\frac{x}{\mu}>0$, $1-\beta-\mu>0$, and $1-\sigma(1-\mu)<0$, we know that (29) is positive, i.e., the LHS of (28) is strictly increasing in $\mu$. Thus, if a solution $\mu$ to (28) exists, it is unique. The solution exists when the LHS of (28) is smaller than the right-hand side, $x$, as $\mu$ goes to 0 and is larger than $x$ as $\mu$ goes to $1-\beta$. Computing the limits of the LHS, we have
@@
Thus, there always exists a unique solution to (28).
Case 1: $\alpha>0$.
$\sigma\leq 1$ $\sigma>1$
Case 1a. $1-\beta\sigma>0$ and $\beta\in(0,1)$ $\infty 0<x<0<x<\infty$
    $1-\beta\sigma>0$ and $\beta=0$ $0<x\leq 1$ $0<x\leq 1$
Case 1b. $1-\beta\sigma<0$ and $\beta\in(0,1)$ n.a. $0<x<\infty$
    $1-\beta\sigma<0$ and $\beta=1$ $0<x<\alpha(\sigma-1)$
Case 2: $\alpha<0$.
$\sigma\leq 1$ $\sigma>1$
Case 2b. $1-\beta\sigma<0$ and $\beta\in(0,1)$ n.a. $0<x<\infty$
Notes: We let $\underline{x}\equiv\overline{\mu}\left[\frac{1-\beta\sigma}{1-\sigma(1-\overline{\mu})}\right]^{\beta-\frac{1}{\sigma}}\left[\frac{1}{\alpha}\frac{1-\beta-\overline{\mu}}{1-\sigma(1-\overline{\mu})}\right]^{-\beta}$, where $\overline{\mu}\equiv\frac{(1-\beta)\sigma+\Delta}{(1-\beta)\sigma+1}$. The case with $\alpha=0$ corresponds to the CES and requires that $\beta\in(0,1)$. The case with $1-\beta\sigma=0$ also corresponds to the CES and requires that $\sigma>1$. If $\alpha=0$ or $\beta=\frac{1}{\sigma}$, then the markup is given by $\frac{1}{\mu}=\frac{1}{1-\beta}\in(1,\infty)$ or $\frac{1}{\mu}=\frac{\sigma}{\sigma-1}\in(1,\infty)$, respectively, and the admissible range of $x$ is given by $0<x<\infty$. Since these two cases are well known, we omit them from the table.

### v2

@@ Appendix B.1 Proof of Proposition 3
Second, assume that $\sigma=1$. In this case, $\beta\neq 1$ must hold because of the assumption that $1-\beta\sigma>0$. Thus, $\Phi(q)>0$ implies $q<-\frac{\beta}{2\alpha}$. We also know from Table B.1 that $0\leq q<-\frac{\beta}{\alpha}$ must hold. Since $\frac{\beta}{2\alpha}<-\frac{\beta}{\alpha}$ holds, the second-order condition holds for all $0\leq q<-\frac{\beta}{2\alpha}$ but it does not hold for $\frac{\beta}{2\alpha}\leq q<-\frac{\beta}{\alpha}$. Since the range of $q$ is modified, we need to modify that for $\mu$. Noting that $\frac{\partial\mu}{\partial q}=\frac{\partial}{\partial q}\bigl(1-\frac{\alpha q+\beta}{\alpha\sigma q+1}\bigr)=-\frac{\alpha(1-\beta\sigma)}{(\alpha\sigma q+1)^{2}}>0$, we evaluate $\mu=1-\frac{\alpha q+\beta}{\alpha\sigma q+1}$ at $q=0$ and $q=-\frac{\beta}{2\alpha}$ to obtain $1-\beta\leq\mu<\frac{2(1-\beta)}{2-\beta}<1$. Hence, we obtain the range of markups $1<\frac{2-\beta}{2(1-\beta)}<\frac{1}{\mu}\leq\frac{1}{1-\beta}$ for $\sigma=1$.
@@
Last, assume that $\sigma>1$. In this case, $\beta\neq 1$ must hold because of the assumption that $1-\beta\sigma>0$. Thus, $\Phi(q)=0$ has two real roots (B.2). Both roots are positive since $1-\beta+\Delta>0$ and $1-\beta-\Delta>0$. Since $\Phi$ is convex, it is negative between its roots. Since $\frac{1-\beta-\Delta}{\alpha(\sigma-1)}<-\frac{\beta}{\alpha}<-\frac{1-\beta+\Delta}{\alpha(\sigma-1)}$, the second-order condition holds for all $0\leq q<-\frac{1-\beta-\Delta}{\alpha(\sigma-1)}$ but does not hold for $\frac{1-\beta-\Delta}{\alpha(\sigma-1)}\leq q<-\frac{\beta}{\alpha}$. Since the bounds for $q$ are modified, we need to modify those for $\mu$. Noting that $\frac{\partial\mu}{\partial q}=\frac{\partial}{\partial q}\bigl(1-\frac{\alpha q+\beta}{\alpha\sigma q+1}\bigr)=-\frac{\alpha(1-\beta\sigma)}{(\alpha\sigma q+1)^{2}}>0$, we evaluate $\mu=1-\frac{\alpha q+\beta}{\alpha\sigma q+1}$ at $q=0$ and $q=-\frac{1-\beta-\Delta}{\alpha(\sigma-1)}<-\frac{\beta}{\alpha}$ to obtain $1-\beta\leq\mu<\frac{(1-\beta)\sigma+\Delta}{1+(1-\beta)\sigma}<1$. Hence, we obtain the bounds for markups $1<\frac{1+(1-\beta)\sigma}{(1-\beta)\sigma+\Delta}<\frac{1}{\mu}\leq\frac{1}{1-\beta}$ for $\sigma>1$.
@@
Since $\beta\neq 1$ and $1-\beta\sigma<0$, $\Phi(q)=0$ has no real roots. Since $\sigma>1$ and $\beta\not\in\{0,1\}$, $\Phi$ is convex and $\Phi(0)=\beta(1-\beta)>0$, which implies that $\Phi$ is positive for all $q\geq 0$. Hence, the SOC holds for all $q\geq 0$. Therefore, the SOC holds for $0\leq q<-\frac{1-\beta}{\alpha(\sigma-1)}$ and thus for $\frac{1}{1-\beta}\leq\frac{1}{\mu}<\infty$ given in Table B.1 when $\beta\in(0,1)$.
Finally, observe from Table 3 that marginal cost pricing $\frac{1}{\mu}=1$ is admissible only when $\beta=0$ in Case 1a. Note that $\frac{\sigma}{\sigma-1}>1$ holds in Case 1b, $\frac{(1-\beta)\sigma+1}{(1-\beta)\sigma+\Delta}>1$ holds in Case 2a, and that $\beta=0$ is not admissible in Case 2b. $\square$
Appendix B.2 Proof of Proposition 4
@@
$\mu\left[\frac{1-\beta\sigma}{1-\sigma(1-\mu)}\right]^{\beta-\frac{1}{\sigma}}\left[\frac{1}{\alpha}\frac{1-\beta-\mu}{1-\sigma(1-\mu)}\right]^{-\beta}=x.$ (B.3)
Consider the subcase with $\beta\in(0,1]$. To obtain the bottom-right expression of Table 4 from (B.3), we first assume that $\sigma\neq 0$. In that case, we can rewrite (B.3) as follows:
@@
When $\sigma$ goes to 0, equation (B.3) reduces to $\mu{\rm e}^{-(1-\beta-\mu)}\left(\frac{1-\beta-\mu}{\alpha}\right)^{-\beta}=x,$ which can be rewritten as
@@
Consider the subcase with $\beta=0$. Then, (B.3) implies $\mu[\frac{1}{1-\sigma(1-\mu)}]^{-\frac{1}{\sigma}}=x$. We can rewrite it as $\frac{1}{1-\sigma(1-\mu)}=(\frac{x}{\mu})^{-\sigma}$, so that $1-\sigma(1-\mu)=(\frac{\mu}{x})^{-\sigma}$. Hence, we obtain $1-\mu=\frac{1-(\frac{\mu}{x})^{-\sigma}}{\sigma}$. When $\sigma$ goes to 0, the foregoing equation reduces to $1-\mu=\ln(\frac{\mu}{x})$.
@@
We now prove the existence and uniqueness of a markup function. The right-hand side of expression (B.3) is exogenous for each firm, whereas the left-hand side depends on the markup $\mu$ that is endogenously determined. Differentiating the left-hand side of (B.3) with respect to $\mu$ yields
$\frac{\partial{\rm LHS}}{\partial\mu}=\frac{1-\beta-\mu^{2}-\sigma(1-\beta)(1-\mu)^{2}}{(1-\beta-\mu)[1-\sigma(1-\mu)]}\frac{x}{\mu}.$ (B.4)
@@
$\frac{1-\beta-\mu^{2}-\sigma(1-\beta)(1-\mu)^{2}}{(1-\mu)(1-\beta\sigma)}>0,$ (B.5)
@@
Consider the subcase with $\beta\in(0,1)$. In this case, $q>0$ must hold as explained before, so that $1-\beta-\mu>0$. Since $1-\mu>\beta>0$ and $1-\beta\sigma>0$ hold, the denominator of (B.5) is positive. We thus obtain $1-\beta-\mu^{2}-\sigma(1-\beta)(1-\mu)^{2}>0$ from (B.5).
Since $\frac{x}{\mu}>0$, $1-\beta-\mu>0$, and $1-\sigma(1-\mu)>0$, we know that (B.4) is positive, i.e., the LHS of (B.3) is strictly increasing in $\mu$. Thus, if a solution $\mu$ to (B.3) exists, it is unique. The solution exists when the LHS of (B.3) is smaller than the right-hand side, $x$, as $\mu$ goes to 0 and is greater than $x$ as $\mu$ goes to $1-\beta$. Computing the limits of the LHS, we have
@@
Thus, when $\sigma\leq 1$ and $0<\beta<1$ hold, there exists a unique solution $\mu\in(0,1-\beta)$ to (B.3) regardless of the value of $x$.
Consider the subcase with $\beta=0$. Setting $\beta=0$ and differentiating the LHS of (B.3) yield $1-\sigma(1-\mu)]^{-\frac{\sigma-1}{\sigma}}[1-\sigma(1-\mu)+\mu]>0$. Computing the limits of the LHS, we have
@@
Thus, when $\sigma\leq 1$ and $\beta=0$ hold, there exists a unique solution $\mu\in(0,1]$ to (B.3) only for $x\leq 1$.
@@
Consider the subcase with $\beta\in(0,1)$. In this case, $q>0$ must hold, so that $1-\beta-\mu>0$. Since $1-\mu>\beta>0$ and $1-\beta\sigma>0$ hold, the denominator of (B.5) is positive. We thus obtain $1-\beta-\mu^{2}-\sigma(1-\beta)(1-\mu)^{2}>0$ from (B.5).
Since $\frac{x}{\mu}>0$, $1-\beta-\mu>0$, and $1-\sigma(1-\mu)>0$, we know that (B.4) is positive, i.e., the LHS of (B.3) is strictly increasing in $\mu$. Thus, if a solution $\mu$ to (B.3) exists, it is unique. The solution exists when the LHS of (B.3) is smaller than the right-hand side, $x$, as $\mu$ goes to $\frac{\sigma-1}{\sigma}$ and is greater than $x$ as $\mu$ goes to $1-\beta$. Computing the limits of the LHS, we have
@@
Thus, when $\sigma>1$ and $0<\beta<1$ hold, there exists a unique solution $\mu\in\bigl(\frac{\sigma-1}{\sigma},1-\beta\bigr)$ to (B.3) regardless of the value of $x$.
Consider the subcase with $\beta=0$. Setting $\beta=0$ and differentiating the LHS of (B.3) yield $1-\sigma(1-\mu)]^{-\frac{\sigma-1}{\sigma}}[1-\sigma(1-\mu)+\mu]>0$. Computing the limits of the LHS, we have
@@
Consider the subcase with $\beta\in(0,1)$. The range of markups $\frac{\sigma}{\sigma-1}<\frac{1}{\mu}\leq\frac{1}{1-\beta}$ in Table 3 implies that the range of $\mu$ is given by $1-\beta\leq\mu<\frac{\sigma-1}{\sigma}$. Since $\mu<\frac{\sigma-1}{\sigma}<1$ and $1-\beta\sigma<0$ hold, the denominator of (B.5) is negative. We thus obtain $1-\beta-\mu^{2}-\sigma(1-\beta)(1-\mu)^{2}<0$ from (B.5).
Recall that when $\beta\in(0,1)$, $q>0$ must hold, so that $1-\beta-\mu<0$. Since $\frac{x}{\mu}>0$, $1-\beta-\mu<0$, and $1-\sigma(1-\mu)<0$, we know that (B.4) is negative, i.e., the LHS of (B.3) is strictly decreasing in $\mu$. Thus, if a solution $\mu$ to (B.3) exists, it is unique. The solution exists when the LHS of (B.3) is greater than the right-hand side, $x$, as $\mu$ goes to $1-\beta$ and is smaller than $x$ as $\mu$ goes to $\frac{\sigma-1}{\sigma}$. Computing the limits of the LHS, we have
@@
Thus, when $0<\beta<1$, there exists a unique solution $\mu\in\bigl(1-\beta,\frac{\sigma-1}{\sigma}\bigr)$ to (B.3) regardless of the value of $x$.
Consider the subcase with $\beta=1$. The range of markups $\frac{\sigma}{\sigma-1}<\frac{1}{\mu}<\infty$ in Table 3 implies that the range of $\mu$ is given by $0<\mu<\frac{\sigma-1}{\sigma}$. Since $\mu<\frac{\sigma-1}{\sigma}<1$ and $1-\beta\sigma<0$ hold, the denominator of (B.5) is negative. We thus obtain $1-\beta-\mu^{2}-\sigma(1-\beta)(1-\mu)^{2}<0$ from (B.5).
Recall that when $\beta=1$, $q>0$ must hold, so that $1-\beta-\mu<0$. Since $\frac{x}{\mu}>0$, $1-\beta-\mu<0$, and $1-\sigma(1-\mu)<0$, we know that (B.4) is negative, i.e., the LHS of (B.3) is strictly decreasing in $\mu$. Thus, if a solution $\mu$ to (B.3) exists, it is unique. The solution exists when the LHS of (B.3) is greater than the right-hand side, $x$, as $\mu$ goes to 0 and is smaller than $x$ as $\mu$ goes to $\frac{\sigma-1}{\sigma}$. Computing the limits of the LHS, we have
@@
In this case, $\beta=0$ is excluded by Step 1 of the proof of Proposition 3 and $\beta=1$ is not admissible by Step 2 of the proof of Proposition 3. The range of markups $\frac{(1-\beta)\sigma+1}{(1-\beta)\sigma+\Delta}<\frac{1}{\mu}\leq\frac{1}{1-\beta}$ in Table 3 implies that the range of $\mu$ is given by $1-\beta\leq\mu<\frac{(1-\beta)\sigma+\Delta}{(1-\beta)\sigma+1}$. Since $\mu<\frac{(1-\beta)\sigma+\Delta}{(1-\beta)\sigma+1}<1$ and $1-\beta\sigma>0$ hold, the denominator of (B.5) is positive. We thus obtain $1-\beta-\mu^{2}-\sigma(1-\beta)(1-\mu)^{2}>0$ from (B.5).
Recall that when $\beta\in(0,1)$, $q>0$ must hold, so that $1-\beta-\mu<0$. Since $\frac{x}{\mu}>0$, $1-\beta-\mu<0$, and $1-\sigma(1-\mu)>0$, we know that (B.4) is negative, i.e., the LHS of (B.3) is strictly decreasing in $\mu$. Thus, if a solution $\mu$ to (B.3) exists, it is unique. The solution exists when the LHS of (B.3) is greater than the right-hand side, $x$, as $\mu$ goes to $1-\beta$ and is smaller than $x$ as $\mu$ goes to $\overline{\mu}\equiv\frac{(1-\beta)\sigma+\Delta}{(1-\beta)\sigma+1}$. Computing the limits of the LHS, we have
@@
Thus, there exists a unique solution $\mu\in(1-\beta,\overline{\mu})$ to (B.3) for $x>\underline{x}$.
@@
In this case, $\beta=0$ and $\beta=1$ are not admissible, whereas $\sigma>1$ must hold. The range of markups $\frac{1}{1-\beta}\leq\frac{1}{\mu}<\infty$ in Table 3 implies that the range of $\mu$ is given by $0<\mu\leq 1-\beta$. Since $\mu\leq 1-\beta<1$ and $1-\beta\sigma<0$ hold, the denominator of (B.5) is negative. We thus obtain $1-\beta-\mu^{2}-\sigma(1-\beta)(1-\mu)^{2}<0$ from (B.5).
Recall that when $\beta\in(0,1)$, $q>0$ must hold, so that $1-\beta-\mu>0$. Since $\frac{x}{\mu}>0$, $1-\beta-\mu>0$, and $1-\sigma(1-\mu)<0$, we know that (B.4) is positive, i.e., the LHS of (B.3) is strictly increasing in $\mu$. Thus, if a solution $\mu$ to (B.3) exists, it is unique. The solution exists when the LHS of (B.3) is smaller than the right-hand side, $x$, as $\mu$ goes to 0 and is larger than $x$ as $\mu$ goes to $1-\beta$. Computing the limits of the LHS, we have
@@
Thus, there always exists a unique solution to (B.3).